# Attention Projection Mixing and Exogenous Anchors

### v1

@@ 4.3 Extending Mixing to Q, K, and G Pathways
In contrast to $Q$ and $K$, residual connections for the gating logits $G$ were straightforward to incorporate even without normalization. We hypothesize that gate logits, which are processed through a sigmoid nonlinearity, are naturally less sensitive to distributional mismatch because the sigmoid compresses their scale.
4.4 Mix-Compress-Refine Theory, Gated Attention, and ResFormer
(a) Attention entropy by layer.
(b) Maximum activations by layer.
(c) PCA Core Features by Layer.
(d) First-token attention (attention sink) by layer.
(e) Average cosine similarity between different token representations within the same layer.
(f) Cosine similarity between the current layer’s representation and the input embedding.
Figure 2: Attention-pattern and representation analysis across model variants. Elementwise is used unless stated otherwise. Some graphs include input embeddings for comparison.
As shown in Figure 2, the baseline Transformer’s behavior aligns with the Mix-Compress-Refine theory proposed by Queipo-de-Llano et al. (2025): (1) it begins with a high attention entropy phase for broad, contextual integration of token information, (2) transitions into a compression valley marked by dominant attention sinks that halt mixing and reduce representational dimensionality to filter out useless information, (3) concludes with a sudden rise in attention entropy as sinks dissipate, enabling the refined, token-specific processing necessary for generation.
While these three stages emerge naturally in standard Transformers, they are not optimized for computational efficiency. The model first spends valuable layers collapsing contextual information and then spends more layers slowly building it back up during stage 2 (Figure 2(c)).
The gating mechanism appears to address the model’s need to filter irrelevant context, a core function of stage 2. By introducing input-dependent sparsity that selectively modulates information flow at every layer, gated attention seems to provide a form of distributed filtering. This may reduce or eliminate the need for a sharp, dedicated compression phase characterized by dominant attention sinks. The empirical evidence is consistent with this interpretation: as shown in Figure 2(c) and Table 2, models with gated attention exhibit no clear compression valley and show a drastic reduction in attention sink magnitude compared to the baseline. This suggests the gates could act as a learned, adaptive replacement for the more rigid, sink-based filtering.
Metric (Avg) Baseline Dynamic ExoFormer ExoFormer Gated NuResFormer ResFormer
Attention Sink 0.2572 0.0073 0.0041 0.0091 0.0112 0.0212
Token Similarity 0.5365 0.8622 0.8265 0.6328 0.5783 0.6558
PCA Core Features 838.9 752.1 818.3 880.0 887.9 853.8
Table 2: ExoFormer and its dynamic variant exhibit the lowest intrinsic dimensionality (PCA Core Features) and high token similarity, indicative of representation collapse, yet achieve the lowest perplexity and highest downstream accuracy. ExoFormer further eliminates attention sink compared to standard Gated Attention.
4.4.1 The Offloading Hypothesis: Specialization via Exogenous Anchors
We quantify representation collapse using two complementary metrics.
  1. 1.
    Token Similarity: Figure 2(e) plots the average cosine similarity between different token representations within the same layer. While all models exhibit an increase in similarity with depth, ExoFormer is an extreme case, with layer token similarity peaking at 95% and 93% for the dynamic variant. This indicates that, internally, token representations become nearly indistinguishable.
  2. 2.
    Intrinsic Dimensionality: As shown in Figure 2(c), most models maintain a high intrinsic dimensionality. In stark contrast, ExoFormer’s core features remain consistently below other models, with the final layer utilizing only approximately 738 core features (716 for the dynamic variant), compared to 835 for the baseline.
We note that residual normalizations are unlikely to be responsible for this over-smoothing, given that NuResFormer, employing full residual norms, yields the greatest mean of PCA core features, while E-ExoFormer (no norm) displays similar signs of collapse.
Furthermore, analyzing the trajectory of hidden states reveals a striking divergence from the starting point. Figure 2(f) shows the cosine similarity between each layer’s representation and that of the input embedding. In ExoFormer, this similarity starts extremely low and steadily decays, with deeper layers becoming nearly orthogonal or even slightly anti-correlated with the initial representation.
The ”Offloading” Hypothesis
Standard theory posits that over-smoothing degrades model performance, as it implies a loss of information necessary for prediction. The fact that dynamic E-ExoFormer achieves the highest average accuracy (50.27%) and lowest perplexity (14.09) while exhibiting internal smoothing—near-identical token representations and low intrinsic dimensionality—presents a clear contradiction.
To resolve this paradox, we propose the Offloading Hypothesis: the exogenous anchor allows the sequential layers to specialize almost exclusively in the final ”refinement” stage (Stage 3) of the Mix-Compress-Refine cycle. Evidence for this specialization is visible in the token-similarity trajectory (Figure 2(e)). All models show a pattern where similarity rises to an early maximum, dips to a local minimum, rises again, and finally falls in the last layer. We interpret the local minimum as the start of Stage 3: the refinement phase dedicated to preparing representations for next-token prediction. Notably, ExoFormer variants spend approximately two-thirds of their layers in this stage, the highest proportion of any model, while the baseline spends only one-third.
This specialization is possible because exogenous anchors fundamentally change how information is managed. In a standard Transformer, the residual stream must concurrently carry two critical types of information: (1) distinct features that preserve token identity (”What token am I?”), and (2) transformed, task-relevant features that evolve through the layers to support next-token prediction. Over-smoothing in a standard model catastrophically loses the first type, crippling the model’s ability to route information. ExoFormer circumvents this by introducing a permanent reference library, the exogenous anchor, that is re-injected at every layer.
  1. 1.
    The Anchor assumes the ”Identity” role. By providing a guaranteed, high-fidelity source of token distinctiveness at every layer, the exogenous anchor relieves the main residual stream from the duty of preserving static token features. The dedicated anchor parameters ($W^{Q}_{\text{anc}},W^{K}_{\text{anc}},W^{V}_{\text{anc}},W^{G}_{\text{anc}}$) can specialize exclusively in maximizing variance and distinctiveness.
  2. 2.
    The Sequential Layers specialize in ”Computation.” With identity information externally supplied, the internal layers are liberated from the constraint of preserving early-layer geometry. The observed ”collapse” can thus be reinterpreted as a gradual signal purification. The layers act as aggressive filters, compressing the representation to focus solely on the progressive, step-by-step computation required for reasoning, while dynamically discarding redundant or static information.
This clean separation allows each component to excel at its specialized function, directly causing the observed extremes: maximal internal smoothing and minimal similarity to the first layer. Further analysis is warranted to fully characterize the dynamics of this smoothing effect and its precise relationship to performance, but the evidence suggests that ExoFormer’s ”collapse” is a feature of its architecture.
5 Conclusion
We introduce NuResFormer, a framework enabling normalized cross-layer residual mixing across all attention pathways, and identify the architectural tension when the first layer must serve as both a reusable anchor and a progressive computational block. To resolve this, we propose ExoFormer, which learns dedicated exogenous anchor projections, decoupling these roles and consistently outperforming NuResFormer variants. We connect these improvements to the Mix-Compress-Refine theory and propose the offloading hypothesis: external anchors preserve token identity, allowing internal layers to specialize in refinement, which explains ExoFormer’s superior efficiency and performance despite signs of internal representation collapse.
6 Limitations
The experiments were conducted at a modest scale (450M parameters, 10B tokens), and the observed behaviors may not generalize to larger models or longer training. Due to resource constraints, we did not perform an exhaustive search for the optimal initialization of the coefficient $\lambda$. The evaluation focuses on multiple-choice benchmarks, leaving the impact on long-context tasks unknown. Finally, while the results are empirically grounded, a formal theoretical explanation for why normalized reuse works is still lacking.
Appendix A Complete Evaluation Results
Model ARC-c ARC-e HellaSwag OBQA PIQA Wino Avg. Acc PPL Params (M)
Baselines
Base Transformer 30.97 63.38 42.02 33.60 67.30 51.54 48.14 14.79 454
Gated Attention 30.89 64.69 42.73 34.00 67.14 53.35 48.80 14.64 453
ResFormer (Value Residual) 33.62 64.86 43.49 34.40 68.88 52.64 49.65 14.32 454
Naïve Combination (ResFormer+Gate) 32.94 64.52 43.47 32.60 68.34 52.64 49.09 14.25 453
NuResFormer Variants (Internal Anchor)
E-NuResFormer (only Q/K residual norms) 31.74 64.27 43.83 33.00 68.23 52.41 48.91 14.21 453
H-NuResFormer (only Q/K residual norms) 34.73 64.06 44.45 34.20 67.57 53.04 49.68 14.21 453
S-NuResFormer (only Q/K residual norms) 32.85 66.08 43.87 34.20 68.28 53.67 49.83 14.22 453
E-NuResFormer (no norm, {V,G} only) 32.42 64.94 43.80 35.00 67.52 51.54 49.20 14.24 453
H-NuResFormer (no norm, {V,G} only) 31.23 64.18 43.50 34.00 67.68 54.06 49.11 14.22 453
S-NuResFormer (no norm, {V,G} only) 31.83 64.44 43.66 34.80 68.44 52.49 49.28 14.24 453
E-NuResFormer (full) 33.70 65.07 44.23 33.60 68.34 53.12 49.68 14.15 453
H-NuResFormer (full) 33.79 65.11 44.09 33.40 67.41 52.72 49.42 14.17 453
S-NuResFormer (full) 32.42 64.02 43.93 33.20 67.90 53.20 49.11 14.17 453
ExoFormer Variants (External Anchor)
E-ExoFormer (no norm) 32.08 63.76 43.49 34.40 68.77 55.64 49.69 14.30 457
Dynamic E-ExoFormer 33.36 65.87 44.54 34.40 68.28 55.17 50.27 14.09 457
E-ExoFormer (full) 34.73 64.65 44.28 36.40 67.74 53.59 49.85 14.13 457
H-ExoFormer (full) 32.17 65.87 44.00 32.40 68.23 54.06 49.23 14.14 457
S-ExoFormer (full) 32.17 65.66 44.33 33.60 68.06 51.14 49.16 14.15 457
Table 3: Comprehensive evaluation results on 6 multiple-choice downstream tasks (accuracy in %), validation perplexity (PPL), and parameter counts. Dynamic E-ExoFormer achieves the highest overall accuracy and lowest perplexity. Full residual norms are utilized unless specified otherwise. ”no norm” indicates no residual normalization was applied to the anchor signal. Bold values indicate the best performance for each metric.
Appendix B Residual Mixing Granularity and Emergent Head-level Structure
A central design choice in our unified residual framework is the granularity at which mixing coefficients $\lambda$ are learned. We evaluate three levels—scalar (S), headwise (H), and elementwise (E)—across both NuResFormer and ExoFormer variants.
@@ Appendix B Residual Mixing Granularity and Emergent Head-level Structure
Figure 3: Heatmaps showing the learned mixing coefficient ratio $\lambda_{n,1}/\lambda_{n,2}$ for each residualized component $\{Q,K,V,G\}$ across layers (y-axis) and channels/heads (x-axis). This ratio quantifies the model’s reliance on the anchor relative to the current layer’s projection; a higher value indicates stronger reuse of the early signal. Even with elementwise freedom, the learned pattern is largely constant within heads for $V$, while $Q$, $K$, and $G$ exhibit finer-grained structure. The dynamic variant shows more uniform distributions across channels and layers.
Emergent Head Structure in Elementwise Mixing.
A striking qualitative finding emerges from visualizing the learned elementwise coefficients, as shown in Figure 3. For the value pathway ($V$), the heatmaps exhibit sharp boundaries that align precisely with head blocks. Within each head—a contiguous span of $d_{k}$ channels—the mixing coefficients $\lambda^{V}_{n,1}$ and $\lambda^{V}_{n,2}$ are nearly constant, with primary discontinuities occurring between heads. This structure emerges despite the optimizer having the freedom to set each channel independently. The result implies that the natural, discovered unit of residual control for content ($V$) is the attention head as a whole, aligning with previous research on heads as specialized submodules for routing information (Voita et al., 2019).
The patterns for queries ($Q$), keys ($K$), and gate logits ($G$) are markedly different. These components frequently exhibit finer-grained, intra-head structure, such as alternating bands of high and low reuse (“striping”). This suggests a form of sub-head specialization, where specific channel subspaces within a head are selectively reused or suppressed. Notably, the positions of these high/low bands in $Q$ often align with those in $K$.
Flexibility Enabled by Dynamic Mixing.
As shown in Figure 3(d), $\lambda_{n,1}/\lambda_{n,2}$ for queries, keys, and gating logits are substantially more uniformly distributed across channels and layers in the dynamic variant, with fewer banded structures seen in static mixing. Rather than converging to a fixed, layer-specific reuse policy, the model can modulate the strength of the anchor pathway for each component in real time based on the input sequence. This flexibility is particularly beneficial for the routing ($Q$, $K$) and selection ($G$) components, whose optimal mixing strategy may depend heavily on immediate contextual needs.
Appendix C Mixing Coefficient
Figure 4: The ratio $\lambda_{n,1}/\lambda_{n,2}$ plotted for each component ($Q$, $K$, $V$, $G$) across layers for models using elementwise mixing. Values greater than 1 indicate stronger reliance on the anchor signal, while values less than 1 indicate preference for current-layer projections.
Figure 4 shows the layer-wise evolution of the mixing coefficient ratio $\lambda_{n,1}/\lambda_{n,2}$ for each attention component.
Appendix D Gate Activation Profiles and First-Layer Selectivity
@@
Figure 5: Depth-wise analysis of gating behavior across model variants.
@@
ExoFormer exhibits an attenuated version of the same profile (Figure 5(a)); its first-layer activation is elevated compared to standalone gating but lower than NuResFormer’s. As shown in Figure 5(b), offloading the anchor role allows ExoFormer to partially restore the first layer’s capacity for selective gating, achieving intermediate sparsity levels between NuResFormer and standalone Gated Attention.
Appendix E Hidden State Similarity
@@
Figure 6: Pairwise cosine similarity of hidden states across transformer layers for models using elementwise mixing. Brighter colors indicate higher similarity.
Appendix F Hyperparameters

### v2

@@ 4.3 Extending Mixing to Q, K, and G Pathways
In contrast to $Q$ and $K$, residual connections for the gating logits $G$ were straightforward to incorporate even without normalization. We hypothesize that gate logits, which are compressed by the sigmoid function, are naturally less sensitive to distributional mismatch.
@@
Emergent Head Structure in Elementwise Mixing.
Visualizing the learned elementwise coefficients reveals emergent patterns specific to each attention component (Figure 3). For the value pathway ($V$), the heatmaps exhibit sharp boundaries that align precisely with head blocks. This structure emerges despite the optimizer having the freedom to set each channel independently, aligning with previous research on heads as specialized submodules for routing information (Voita et al., 2019).
Conversely, queries ($Q$), keys ($K$), and gate logits ($G$) exhibit finer-grained, intra-head structure, such as alternating bands of high and low reuse (“striping”), suggesting a form of sub-head specialization. Notably, the positions of these high/low bands in $Q$ often align with those in $K$.
Flexibility Enabled by Dynamic Mixing.
As shown in Figure 3(a), $\lambda_{n,1}/\lambda_{n,2}$ for queries, keys, and gating logits are substantially more uniformly distributed across channels and layers in the dynamic variant. Rather than converging to a fixed, layer-specific reuse policy, the model can modulate the strength of the anchor pathway for each component in real time based on the input sequence, enabling more freedom in expression in the mixing coefficients themselves.
4.4 Mix-Compress-Refine Theory, Gated Attention, and ResFormer
(a) Attention entropy by layer.
(b) Average cosine similarity between different token representations within the same layer.
(c) PCA Core Features by layer to explain 99% variance.
(d) First-token attention (attention sink) by layer.
Figure 4: Attention-pattern and representation analysis across model variants $\sim\!450M$. Elementwise is used unless stated otherwise. Some graphs include input embeddings for comparison.
As shown in Figure 4, the baseline Transformer’s behavior aligns with the Mix-Compress-Refine theory proposed by Queipo-de-Llano et al. (2025): (1) it begins with a high attention entropy phase for broad, contextual integration of token information, (2) transitions into a compression valley marked by dominant attention sinks and a drastic drop in core features that halt mixing and reduce representational dimensionality to filter out useless information, (3) concludes with a sudden rise in attention entropy as sinks dissipate, enabling the refined, token-specific processing necessary for generation.
While these three stages emerge naturally in standard Transformers, they are not optimized for computational efficiency. The model expends computational capacity collapsing contextual information, only to subsequently reconstruct it, indicating inefficiency in the standard architecture (Figure 4(c)).
The gating mechanism addresses the model’s need to filter irrelevant context, a core function of stage 2. By introducing input-dependent sparsity that selectively modulates information flow at every layer, gated attention provides a form of distributed filtering. This may reduce or eliminate the need for a sharp, dedicated compression phase characterized by dominant attention sinks. The empirical evidence is consistent with this interpretation: as shown in Figure 4(c) and Table 2, models with gated attention exhibit no clear compression valley and show a drastic reduction in attention sink magnitude compared to the baseline. This suggests the gates could act as a learned, adaptive replacement for the more rigid, sink-based filtering.
These findings suggest that both mechanisms alleviate the abrupt information bottleneck characteristic of stage 2, replacing inefficient, sink-based filtering with a more continuous and learned process for isolating relevant information.
4.5 The Offloading Hypothesis: Specialization via Exogenous Anchors
Table 2: Metrics averaged across layers for $\sim\!450M$ models.
Model Attention Sink Token Similarity PCA Core Features
Baseline 0.2580 0.5364 839.0
Dynamic ExoFormer 0.0453 0.6606 885.0
ExoFormer 0.0078 0.5682 890.1
Gated 0.0091 0.6327 880.0
NuResFormer 0.0077 0.5785 887.9
ResFormer 0.0212 0.6558 853.8
The architectural decoupling in ExoFormer enables an interesting functional specialization, which we formalize as the Offloading Hypothesis. By providing a dedicated, high-fidelity source of token identity, the exogenous anchor allows the sequential layers to offload the burden of preserving static features and specialize almost exclusively in the final ”refinement” stage (Stage 3) of the Mix-Compress-Refine cycle.
Evidence for this specialization is clearly visible in the token-similarity trajectory (Figure 4(b)). All models exhibit a characteristic pattern: similarity rises to an early maximum (Stage 1: mixing), dips to a local minimum (Stage 2: compression), rises again, and finally falls in the last layer (Stage 3: refinement). We interpret the local minimum as the onset of Stage 3—the refinement phase dedicated to preparing discriminative representations for next-token prediction. Crucially, ExoFormer variants and NuResFormer spend approximately two-thirds of their layers in this specialized refinement stage, the highest proportion of any model, while the baseline spends only one-third.
This specialization is possible because exogenous anchors fundamentally reconfigure information management within the Transformer. In a standard architecture, the residual stream must concurrently carry two conflicting types of information: (1) distinct, static features that preserve token identity (”What token am I?”), and (2) transformed, task-relevant features that evolve through the layers to support next-token prediction. Over-smoothing in standard models catastrophically loses the first type, crippling the model’s ability to route information effectively. ExoFormer circumvents this tension by introducing a permanent reference library—the exogenous anchor—that is reinjected at every layer, guaranteeing access to high-fidelity token identity.
  1. 1.
    The Anchor Specializes in ”Identity Preservation.” By providing a guaranteed, high-fidelity source of token distinctiveness at every layer, the exogenous anchor relieves the main residual stream from the duty of preserving static token features. The dedicated anchor parameters ($W^{Q}_{\text{anc}},W^{K}_{\text{anc}},W^{V}_{\text{anc}},W^{G}_{\text{anc}}$) can specialize exclusively in maximizing variance and distinctiveness, optimized to serve as a universal reference.
  2. 2.
    The Sequential Layers Specialize in ”Progressive Computation.” With identity information externally supplied, the internal layers are liberated from the constraint of preserving early-layer geometry. They are no longer required to maintain a stable, high-fidelity record of token identity within the residual stream—a task that inherently conflicts with the need to transform and abstract features.
Unburdened by the need to preserve static token features, the layers can more efficiently allocate their capacity to learning complex, high‑level patterns, leading to the observed improvements in both perplexity and downstream accuracy. The architecture effectively separates the stable, memory‑like function of identity preservation (handled by the exogenous anchor) from the adaptive, computational function of progressive refinement (handled by the sequential layers), resulting in a more efficient and performant model.
In ablations where we remove the exogenous anchor during inference, the model suffers a catastrophic failure: core features (measured by PCA dimensions required for 99% variance) plummet drastically and reach 321 in the final layer, while token-to-token similarity peaks at 93%. This dramatic collapse demonstrates that the sequential layers, when deprived of their dedicated identity source, lose most discriminative power. The layers were optimized under the specific assumption that identity preservation was offloaded to the anchor. When that anchor is removed, they have no mechanism to maintain token distinctiveness.
5 Conclusion
We introduced ExoFormer, a novel Transformer architecture that decouples token identity preservation from computational refinement by learning dedicated exogenous anchor projections. Through a unified normalized mixing framework across queries, keys, values, and gates, ExoFormer resolves the architectural tension inherent in reusing first-layer projections. Experiments demonstrate that this decoupling consistently improves performance, with the dynamic variant achieving significant gains in downstream accuracy and data efficiency compared to standard gated attention baselines. These results validate the Offloading Hypothesis, suggesting that externalizing identity preservation enables sequential layers to specialize exclusively in high-level feature transformation, providing an efficient architectural path for enhancing large language models.
Impact Statement
The primary positive societal impact of this work is environmental sustainability. By reducing the number of tokens required to train high-performing models, ExoFormer directly addresses the escalating energy consumption and carbon footprint associated with training large language models (LLMs). The open-source release of code and pre-trained models promotes democratization, lowering the computational barriers for researchers and smaller organizations to develop and deploy state-of-the-art NLP tools. Furthermore, by providing a theoretical framework for how information flows through deep networks (the Offloading Hypothesis), this work aids in the interpretability and mechanistic understanding of LLMs, which is crucial for diagnosing failure modes and ensuring model reliability.
While our contribution is architectural—focused on efficiency and structure rather than specific content generation—it facilitates the deployment of more capable AI systems that could be employed maliciously. Additionally, while our ablations suggest robustness, there is inherent uncertainty in how the ”offloading” mechanism will scale to significantly larger parameter counts (e.g., 100B+), which may introduce novel failure modes or safety challenges not observed in the 450M–1B range studied here.
Appendix A Anchor Reliance and Hidden State Similarity
@@
Figure 5: Heatmaps showing the learned mixing coefficient ratio $\lambda_{n,1}/\lambda_{n,2}$ for each residualized component $\{Q,K,V,G\}$ across layers (y-axis) and channels/heads (x-axis) for model variants $\sim\!450M$. This ratio quantifies the model’s reliance on the anchor relative to the current layer’s projection; a higher value indicates stronger reuse of the early signal. Even with elementwise freedom, the learned pattern is largely constant within heads for $V$, while $Q$, $K$, and $G$ exhibit finer-grained structure. The dynamic variant shows more uniform distributions across channels and layers.
Figure 6: The ratio $\lambda_{n,1}/\lambda_{n,2}$ plotted for each component ($Q$, $K$, $V$, $G$) across layers for models using elementwise mixing for model variants $\sim\!450M$. Values greater than 1 indicate stronger reliance on the anchor signal, while values less than 1 indicate preference for current-layer projections.
Appendix B Gate Activation Profiles and First-Layer Selectivity
@@
Figure 7: Analysis of gating behavior across model variants $\sim\!450M$.
@@
ExoFormer exhibits an attenuated version of the same profile (Figure 7(a)); its first-layer activation is elevated compared to standalone gating but lower than NuResFormer’s. As shown in Figure 7(b), offloading the anchor role allows ExoFormer to partially restore the first layer’s capacity for selective gating.
@@
Figure 8: Pairwise cosine similarity of hidden states across transformer layers for models using elementwise mixing $\sim\!450M$. Brighter colors indicate higher similarity.
Appendix C Complexity Analysis
We present a complexity analysis of ExoFormer variants. Let $L$ be the number of Transformer layers and $d$ the model dimension ($d_{\text{model}}$). We omit RMSNorm because it is negligible in terms of parameters and computation.
C.1 Parameter Overhead
We analyze the parameter overhead of ExoFormer variants relative to a baseline Transformer with Gated Attention. Throughout, we ignore input and output embedding parameters, as they are shared across all models.
Baseline Parameters:
The baseline Transformer with Gated Attention has parameters per layer for the attention projections (queries, keys, values, gate logits and output) and the two-layer SwiGLU feed-forward network. The attention and output projections require $5d^{2}$ parameters (five $d\times d$ matrices), and the FFN requires $6d^{2}$ parameters (assuming expansion factor 4). Thus, ignoring the input and output embedding, the total parameters for $L$ layers are:
$P_{\text{base}}=L\cdot\left(5d^{2}+6d^{2}\right)=11Ld^{2}.$ (13)
Exogenous Anchor Parameters:
ExoFormer introduces dedicated projection matrices for the exogenous anchor: four $d\times d$ matrices, one for each attention component (Q, K, V, G):
$\Delta P_{\text{anchor}}=4d^{2}.$ (14)
Static Mixing Parameters:
For static mixing with elementwise granularity, each layer learns two mixing coefficients (one for anchor, one for current projection) for each of the four components. This amounts to $8d$ parameters per layer:
$\Delta P_{\text{static}}=L\cdot 8d.$ (15)
Dynamic Mixing Parameters:
The Dynamic Mixing (DM) module adds a small MLP per layer. With $d_{\text{DM}}=16$ and $d_{\text{out}}=8$, the dominant term is from the first weight matrix ($d\cdot d_{\text{DM}}$). The second weight matrix and biases (totaling 128 parameters) are negligible for typical $d$:
$\Delta P_{\text{DM}}\approx L\cdot(d\cdot d_{\text{DM}}).$ (16)
Total Parameter Overhead:
Static ExoFormer adds:
$\Delta P_{\text{static ExoFormer}}=4d^{2}+8Ld.$ (17)
Dynamic ExoFormer additionally includes the DM module:
$\Delta P_{\text{dynamic ExoFormer}}=4d^{2}+8Ld+Ld\cdot d_{\text{DM}}.$ (18)
Parameter Ratio:
The extra parameter ratio relative to baseline simplifies to:
$R_{P}^{\text{(static)}}$ $=\frac{4d^{2}+8Ld}{11Ld^{2}}=\frac{4}{11L}+\frac{8}{11d},$ (19)
$R_{P}^{\text{(dynamic)}}$ $=\frac{4d^{2}+8Ld+Ld\cdot d_{\text{DM}}}{11Ld^{2}}=\frac{4}{11L}+\frac{8+d_{\text{DM}}}{11d}.$ (20)
For values ($L=32,d=1024,d_{\text{DM}}=16$), this corresponds to approximately 1.2% overhead for static ExoFormer and 1.3% for dynamic ExoFormer—a modest increase given the observed performance gains.
C.2 Computational Overhead (FLOPs)
We estimate the floating-point operations (FLOPs) per token during the forward pass, considering matrix-vector multiplications as the dominant factor ($2d^{2}$ FLOPs per $d\times d$ matrix). We include the cost of elementwise operations for mixing, although they are computationally smaller ($O(d)$) compared to projections ($O(d^{2})$).
Baseline FLOPs:
The baseline usage corresponds to the parameters utilized at every layer. With $P_{\text{base}}=11Ld^{2}$ (excluding embeddings), the computational cost per token is:
$C_{\text{base}}\approx 2\cdot P_{\text{base}}=22Ld^{2}.$ (21)
Exogenous Anchor FLOPs:
The exogenous anchor projections are computed only once per token using the input embeddings, regardless of the network depth. For the four projection matrices ($W^{Q}_{\text{anc}},W^{K}_{\text{anc}},W^{V}_{\text{anc}},W^{G}_{\text{anc}}$), the cost is:
$\Delta C_{\text{anchor}}\approx 2\cdot(4d^{2})=8d^{2}.$ (22)
Crucially, this cost is constant and does not scale with the number of layers $L$.
Dynamic Mixing FLOPs:
The Dynamic Mixing module operates at every layer. The computational cost includes the projection of the MLP ($d\to d_{\text{DM}}$) and the elementwise mixing operations. The mixing involves two multiplications (coefficients $\times$ anchor/current projections) and one addition per element for the four components (Q, K, V, G), totaling $12d$ FLOPs per layer:
$\Delta C_{\text{DM}}\approx L\cdot(2\cdot d\cdot d_{\text{DM}}+12d).$ (23)
Total FLOPs Overhead:
The total computational overhead ratio is:
$R_{\text{FLOPs}}=\frac{\Delta C_{\text{anchor}}+\Delta C_{\text{DM}}}{C_{\text{base}}}=\frac{8d^{2}+L(2dd_{\text{DM}}+12d)}{22Ld^{2}}=\frac{8}{22L}+\frac{d_{\text{DM}}}{11d}+\frac{6}{11d}.$ (24)
For values ($L=32,d=1024,d_{\text{DM}}=16$), the overhead is dominated by the anchor projection ($\approx 1.1\%$), followed by the dynamic module projection ($\approx 0.14\%$), and the elementwise mixing operations ($\approx 0.05\%$), resulting in a total FLOPs increase of approximately $1.33\%$.
C.3 Efficiency-Performance Trade-off
The FLOPs analysis presented above provides a lower-bound estimate of the computational overhead. In practice, the real-world slowdown may be higher (we observe an increase in latency per token of approximately 8–15% for our reference implementation prioritizing modularity and interpretability over kernel fusion, compared to the theoretical  1.3% FLOPs increase).
We emphasize that our primary contribution is a novel architecture that unifies attention projection mixing, not a production-optimized layer ready for deployment. The net performance benefit even in our setting, evidenced by accuracy gains of  1.5 points and data efficiency improvements of  1.5× against Gated Attention, remains positive even when accounting for the measured runtime overhead, suggesting that an optimized kernel implementation would close the gap with the theoretical 1.3% overhead.
Appendix D Hyperparameters
Table 3: Training hyperparameters for models with different configurations for the main $\sim\!450M$ models. Layer depth was reduced for gated variants to maintain comparable parameter counts. All models were trained on the FineWeb-Edu dataset.